OpenStudy (annie96):

Do I have all of them right. Which number is a solution of 3p > 4p + 6? (Points : 1) –4 –5(Answer) –6 –7 Solve for x: 6x-4y=3 x=1/2+2/3y x=1/2+3/2y x=2+3/2y(Answer) x=2+2/3y Solve for y: 3x-6y=9 y=3/2+1/2x (Answer) y=-3/2+1/2x y=3/2x +1/2 y=2/3 - 1/2x

OpenStudy (math&ing001):

Hey there @Annie96 If you solve the first inequality you get -6>p So which of the choices you think verifies that ?

OpenStudy (annie96):

B.

OpenStudy (math&ing001):

But -5 is not inferior to -6, think again.

OpenStudy (annie96):

-4

OpenStudy (math&ing001):



OpenStudy (annie96):

so its going to be -6

OpenStudy (math&ing001):

It's strictly inferior to -6 so it can't be -6.

OpenStudy (annie96):

-7

OpenStudy (math&ing001):

Yess!

OpenStudy (annie96):

Yay!

OpenStudy (math&ing001):

Haha you finally got it :P

OpenStudy (annie96):

yea i know

OpenStudy (math&ing001):

You want help with the others ?

OpenStudy (annie96):

yes thank you.

OpenStudy (math&ing001):

Ok, so for 6x-4y=3 first get the x-term in one side and the rest to the other side.

OpenStudy (annie96):

4y=6x-3

OpenStudy (math&ing001):

I meant like this 6x=4y+3 then divide everything by 6.

OpenStudy (annie96):

4/6y+3/6 2/3+1/2

OpenStudy (math&ing001):

Yeah but it should stay like an equation : 6x=4y-3 6/6 x = 4/6 y - 3/6 When you simplify you get : x = 2/3 y - 1/2

OpenStudy (annie96):

ok

OpenStudy (annie96):

so the answer would be B.

OpenStudy (math&ing001):

It's x= 2/3 y +1/2 little typo on the - sign So it would be A not B. PS: sorry for making too long to respond I was away.

OpenStudy (annie96):

its ok and thank you for your help.